Hoodmemes:

An airplane flies against the wind from A to B in 8 hours. The same airplane returns from B to A, in the same direction as the wind, in 7 hours. Find the ratio of the speed of the airplane (in still air) to the speed of the wind.

Hoodmemes:

I just need a second opinion on how im anwering this question step by step.

JAIDENGYMNASTIC:

according to question , equation when plane moves in opposite direction L/( x - y) = 8


(1) again, equation when , plane moves in same direction . L/( x + y) = 7
(2) divide (1) by (2) (x + y)/( x - y) = 8/7 7(x + y) = 8( x - y) 7x + 7y = 8x - 8y (8 -7)x = (7 + 8)y x = 15y x/y = 15/1

Hoodmemes:

Let: S be the speed of the airplane in still air, r be the speed of the wind and d the distance between A and B.d = 8(S - r) : airplane flies against the wind,d = 7(S + r) : airplane flies in the same direction as the wind,8(S - r) = 7(S + r),S/r = 15?

Hoodmemes:

i mean we basiclly got the same answer.

JAIDENGYMNASTIC:

LOL